Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 28/3/2017

Location : CR21 in R1 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 8161978083
Leader code: 2735

Actions

  • Hall Manager: water system change-over procedure (stands)
  • Status of miceraid5 disk cleanup (stands)
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel (stands)
  • PFranchini - Look at MOM-DC Configuration eLog tab and update if possible - Added "Temperature" and "Pressure"
  • MUchida - investigate moving Tracker waveguides around to cure "hole" (stands)
  • MCourthold - show JBoehm how to use the Picologger for quench data
  • JBoehm - get the West Wall compressors correctly labelled, as various have been re-purposed
  • VBayliss - confirm shifter cover needed for LH2 startup with Steve Boyd
  • DC - neutron monitor remote readout
  • Actions from previous meeting
  • DC update
  • LH2 update
  • List of Hall activities in the shutdown (Colin)
  • List of issues arising from previous user cycle (may overlap with Colin's list)
  • Recurring list (may overlap with Colin's list)
    • Get a spare EMR board (stands)
    • Compressor 2 SSU communication problems (Ajit is aware; possibly a comms cable)
    • Communication dropouts in the FCD helium level monitor (stands)
    • TKU compressor 2 needs servicing/ spares (stands)
    • TKD may need to swap a waveguide to improve dead channel placement (stands)
    • TKU butterfly helium leak tightening (MU, stands)
    • Status of target1 backup machine in R78 (EO/JL, stands)
    • need to emphasise procedures to gain swipe access in the vac area for new shifters (Steve Boyd)(stands but understood)
    • need a spare PCI board for the tracker ioc pc backup machine (stands)
    • MLCR network isolation test still needs to be done (PF, stands)
    • #1878, #1891 Trench water pump control - pressure appears stable; shifter lockout in progress
    • #1881 Water system PVs not being archived - Services/Environment/Computing spreadsheet (stands)
    • #1880 Alarm Handler not reacting to conventional magnet interlock (stands)
    • #1903, #1881 micestore disk full, also needs to be in Nagios warning checks in place but need testing
    • #1881 File Compactor failure - shifters reported only the "previous error" message; should be in Alarm Handler. Move File Compactor output to top of RC screen? Also waiting on Services/Environment/Computing spreadsheet (stands)
    • VME crates in Hall overheating; crate monitoring looking at wrong crate Services/Environment/Computing spreadsheet, GUI needs fixing (stands)
  • ISIS cycle : May 2 - June 2
  • Shift signup now open on CHEESE
    • Shifts start at daily at midnight now, rather than 8 am

AOB:

Minutes

Present AK,CW,CM,HN,PF,JB,MC

New Actions

  • Ajit - No further development of Hanlet SMS system. Need to propose a basic messaging system still to alleviate need for 24Hr monitoring in person.
  • Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers?
  • Macwaters - Re-make dodgy compressor2 remote cable connector
  • Macwaters - Prepare new F70 compressor for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair.
  • Macwaters - Replace old vac sensors and new valve on cryo 1&2
  • Ajit - Decide on new way to implement what was MOM/DC state machine overide function
  • Macwaters - Report on status of FC psu repair

Previous Actions

  • Hall Manager - water system change-over procedure (stands)
  • EOverton - Status of miceraid5 disk cleanup (stands)
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel (stands)
  • PFranchini Look at MOM-DC Configuration eLog tab and update if possible Added "Temperature" and "Pressure"
  • MUchida - investigate moving Tracker waveguides around to cure "hole". Is Ed happy for this to progress? (stands)
  • MCourthold - show JBoehm how to use the Picologger for quench data. Craig can help. (stands)
  • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
  • VBayliss - confirm shifter cover needed for LH2 startup with Steve Boyd
  • DC - neutron monitor remote readout

Other

  • DC update
    • Cyclone cranes installed trench chain hoist for compressors
    • ICS replaced broken controller on roof water system. Now all ok.
    • Neutron monitor webcam installed. Zero reading is due to no Gamma reading now. System is all good and tested.
    • Cryo 1&2 at room temp. Looking to heat to 315K and then normal pump purge procedure. Will then be ready for coldhead hot swap on April 5th
    • Target BPS test with ISIS planned for 20th April
  • LH2 update - unsettled...
  • List of Hall activities in the shutdown (Colin)
  • List of issues arising from previous user cycle (may overlap with Colin's list)
  • Recurring list (the following now to be included in Colin's list from 28th March)
    • Get a spare EMR board (stands)
    • Compressor 2 SSU communication problems (Ajit is aware; possibly a comms cable)
    • Communication dropouts in the FCD helium level monitor (stands)
    • TKU compressor 2 needs servicing/ spares (stands)
    • TKD may need to swap a waveguide to improve dead channel placement (stands)
    • TKU butterfly helium leak tightening (MU, stands)
    • Status of target1 backup machine in R78 (EO/JL, stands)
    • need to emphasise procedures to gain swipe access in the vac area for new shifters (Steve Boyd)(stands but understood)
    • need a spare PCI board for the tracker ioc pc backup machine (stands)
    • MLCR network isolation test still needs to be done (PF, stands)
    • #1878, #1891 Trench water pump control - pressure appears stable; shifter lockout in progress
    • #1881 Water system PVs not being archived - Services/Environment/Computing spreadsheet (stands)
    • #1880 Alarm Handler not reacting to conventional magnet interlock (stands)
    • #1903, #1881 micestore disk full, also needs to be in Nagios warning checks in place but need testing
    • #1881 File Compactor failure - shifters reported only the "previous error" message; should be in Alarm Handler. Move File Compactor output to top of RC screen? Also waiting on Services/Environment/Computing spreadsheet (stands)
    • VME crates in Hall overheating; crate monitoring looking at wrong crate Services/Environment/Computing spreadsheet, GUI needs fixing (stands)
  • ISIS cycle : May 2 - June 2
  • Shift signup now open on CHEESE
    • Shifts start at daily at midnight now, rather than 8 am

AOB:

Updated by MacWaters, Craig over 6 years ago ยท 6 revisions